OpenStudy (anonymous):

solve the equations (I have the answer which is x=5 but need the steps please 3x/(x-2)=(3x+5)/(x-1)

OpenStudy (anonymous):

use cross multiplication

OpenStudy (anonymous):

3x(x-1)=(3x+5)(x-2) 3x^2-3x=3x^2-6x+5x-10 -2x=-10 x=5

OpenStudy (anonymous):

\[3x \div(x-2) = (3x+5)\div(x-1)\] \[3x ^{2} - 3x = 3x ^{2} + 5x - 6x - 10\] \[3x ^{2} -3x = 3x ^{2} -x -10\] \[3x ^{2} - 3x - 3x ^{2} + x = -10\] \[-2x = -10\] \[2x = 10\] \[x = 5\]

OpenStudy (anonymous):

3x(x-1)=(3x+5)(x-2) or 3x(x-1)=3x(x-2)+5(x-2) or 3x(x-1)-3x(x-2)=5x-10 or 3x(x-1-x+2)=5x-10 or 3x =5x-10 or 2x=10 or x=5